given line1 passing through the points (-4,-1) and (-1,0) and line2 passing through the points (5,8) and (0,4) are the lines parallel, perpendicular, or neigher?
Answer by tutorcecilia(2152)

Lines that are parallel have the same slope.
Lines that are perpendicular have slopes that are negative reciprocals of each other.
.
Line #1 = m=(0-(-1))/-1-(-4) [plug the values into the formula for a slope]
m=1/3
.
Line # 2 =m=(4-8)/0-5)=-4/-5=4/5
.
Looking at the slopes of both lines, the lines are neither parallel nor perpendicular.